OpenStudy (anonymous):

John bought 15 cookies and ate 3 of them. He ate ____% of the cookies.

OpenStudy (anonymous):

well what i know is he ate 1/5 of them. so this means he ate 20% of them

OpenStudy (igreen):

To find the percentage divide then multiply by 100. 3 / 15 = 0.2 * 100 = 20%

OpenStudy (igreen):

Welcome to Open Study! You can give medals by clicking 'Best Response'. @dyamondsr4ever

OpenStudy (anonymous):

I hope I answered the question for you

OpenStudy (anonymous):

20%